OpenStudy (clarence):

What is this question asking me to do? Assuming that the sequence {an} defined below is convergent, find its limit. a1=1,an+1=12−1an,n≥1.

ganeshie8 (ganeshie8):

you need to find \(\lim\limits_{n\to\infty}~a_n\)

OpenStudy (clarence):

\[a _{1}=1,a _{n+1}=12-\frac{ 1 }{ a _{n} }, n \ge1\]

OpenStudy (clarence):

How would I go abouts doing that? @ganeshie8

ganeshie8 (ganeshie8):

use below \(\lim\limits_{n\to\infty}~a_{n+1} = \lim\limits_{n\to\infty}~a_n\)

ganeshie8 (ganeshie8):

let \(\lim\limits_{n\to\infty}~a_n =\lim\limits_{n\to\infty}~a_{n+1} =x\)

ganeshie8 (ganeshie8):

so you want to solve \(x\)

ganeshie8 (ganeshie8):

we have \(a _{n+1}=12-\frac{ 1 }{ a _{n} }\) take limit \(\lim\limits_{n\to\infty} a _{n+1}=\lim\limits_{n\to\infty} 12-\frac{ 1 }{ a _{n} }\\~\\x = 12-\frac{1}{x}\) you can solve \(x\)

OpenStudy (clarence):

\[x=6\pm \sqrt{35}\] But how did you manage to get that from the question?

ganeshie8 (ganeshie8):

get what ?

OpenStudy (clarence):

Get what to do :p

ganeshie8 (ganeshie8):

ahh i have already told you :) you need to be a lil more specific haha

OpenStudy (clarence):

How did you know that you had to equate the two limits together to figure out the question and stuff? :P

ganeshie8 (ganeshie8):

okay honestly i didn't invent that trick... i don't think its some thing that occurs easily to anyone... it needs to be shown by others...

OpenStudy (clarence):

Ahh, fair enough then, well thanks for the assistance as always! :)

ganeshie8 (ganeshie8):

np :)